OpenStudy (anonymous):

Find area between two curves: y=√x , .5x , x=9

OpenStudy (anonymous):

Y=.5x

OpenStudy (anonymous):

First curve: \[\int \sqrt{x}\,dx=\frac{2x^{3/2}}{3}\]

OpenStudy (anonymous):

What do you mean with the rest of the stuff on your question? ".5x , x=9"

OpenStudy (anonymous):

Y= .5x and x=9 ...they are all the functions they gave us

OpenStudy (anonymous):

Find the area of the enclosed region by the given curves

OpenStudy (anonymous):

The question is missing the Y thats all.

OpenStudy (anonymous):

Do you know the formula for the area under the line Y=.5x ? If you don't then you really need to study.

OpenStudy (anonymous):

What do u mean

OpenStudy (anonymous):

It should be an easy question given the subject of the question. If you don't know the answer it means you don't grasp basic concepts implied by the question, like area, curves and curve equations.

OpenStudy (anonymous):

The area under y=0.5x is \(\displaystyle\frac{x^2}{4}\). But the two curves cross in a point besides (0, 0).